# Data-Driven Dynamics Modeling of Miniature Robotic Blimps Using Neural ODEs With Parameter Auto-Tuning

### v1

@@ V-C Performance Analysis on Parameter AutoTuning
Figure 7: The scatter plot of three performance indices $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$, demonstrating that the integration of parameter auto-tuning in the first phase contributes to enhanced modeling accuracy in both the first-principle and neural network modules.
Figure 7 shows the scatter plot of $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$. Each data point corresponds to one individual experimental trial. The clusterxing of points primarily lies in the first quadrant, indicating a positive correlation between the performance gain of parameter auto-tuning in the first phase and that in the second phase. Additionally, the intensity of the red color reflects the degree of performance gain achieved by the entire two-phase model training over the first-principle model, with darker shades indicating higher loss reduction. In the scatter plot, darker red dots are mainly distributed in the first quadrant. This observation suggests that the integration of parameter auto-tuning in the first phase contributes to enhanced modeling accuracy in both the first-principle and neural network modules.
@@
Figure 8: An example spiral trajectory for $(\Delta_{\overline{\bm{r}}_{x}},F_{l},F_{r})=(1\text{cm},1.4\text{gf},5.8\text{gf})$. The predicted trajectory of ABNODE closely aligns with the experimental result, showcasing its superior performance compared to BNODE, which lacks physical parameter auto-tuning.
@@
To visually demonstrate the modeling improvement achieved through parameter auto-tuning, Fig. 8 illustrates an example spiraling flight trajectory in three-dimensional space. From this visualization, we observe that the model prediction of ABNODE aligns closely with the experimental trajectory, demonstrating a significant improvement compared to BNODE which lacks parameter auto-tuning.
@@ VI Conclusion
This letter proposed the ABNODE model that consisted of a first-principle module with physical parameter auto-tuning and a neural network module for modeling residual dynamics in robotic blimps. Extensive spiraling experiments were conducted, the results of which demonstrated improved modeling accuracy and generalizability compared to three baseline models including the first-principle, SINDYc, and KNODE.
In future work, we plan to extend the application scope of the ABNODE model to accommodate a wider range of motions executed by robotic blimps. Additionally, we will investigate feedback control designs built upon the ABNODE model for agile and complex flight tasks.

### v2

@@ V-C Performance Analysis on Parameter AutoTuning
Figure 6: Analysis of parameter auto-tuning. (a) The scatter plot of three performance indices $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$, demonstrating that the integration of parameter auto-tuning in the first phase contributes to enhanced modeling accuracy in both the first-principle and neural network modules. (b) An example spiral trajectory for $(\Delta_{\overline{\bm{r}}_{x}},F_{l},F_{r})=(2\text{cm},1.4\text{gf},5.8\text{gf})$. The predicted trajectory of ABNODE closely aligns with the experimental result, showcasing its superior performance compared to BNODE, which lacks physical parameter auto-tuning.
Figure 6 (a) shows the scatter plot of $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$. Each data point corresponds to one individual experimental trial. The clusterxing of points primarily lies in the first quadrant, indicating a positive correlation between the performance gain of parameter auto-tuning in the first phase and that in the second phase. Additionally, the intensity of the red color reflects the degree of performance gain achieved by the entire two-phase model training over the first-principle model, with darker shades indicating higher loss reduction. In the scatter plot, darker red dots are mainly distributed in the first quadrant. This observation suggests that the integration of parameter auto-tuning in the first phase contributes to enhanced modeling accuracy in both the first-principle and neural network modules.
@@
To visually demonstrate the modeling improvement achieved through parameter auto-tuning, Fig. 6 (b) illustrates an example spiraling flight trajectory in three-dimensional space. From this visualization, we observe that the model prediction of ABNODE aligns closely with the experimental trajectory, demonstrating a significant improvement compared to BNODE which lacks parameter auto-tuning.
V-D Discussions on ABNODE Models
V-D1 Adaptability to prediction horizons and time steps
All models were trained with a prediction time step of 1/60 s in aforementioned experiments. To further investigate, we conducted experiments and evaluated the ABNODE models on a test dataset with sequence lengths exceeding 45 seconds. These experiments employed varying prediction horizons and time steps. Specifically, we evaluated 10 ABNODE models with time steps set at 2/60 s, 3/60 s, 4/60 s, 5/60 s, and prediction horizons set at 30 s, 35 s, 40 s, 45 s. Consequently, varing prediction time steps resulted in different data lengths under consistent total time lengths. Additionally, we conducted experiments with mixed prediction time steps of 1/60 s, 2/60 s, and 3/60 s in a cyclical pattern. Table III shows the average percentage change in loss compared to a time step of 1/60 s across varying time steps under the same corresponding prediction horizons. We observe that while different prediction time steps do result in variations in prediction loss, the average loss increase remains below 0.4% across all selected prediction horizons, indicating that the ABNODE model has satisfactory adaptability to varying time steps.
TABLE III: Average loss percent change of ABNODE models under various time steps compared to the 1/60 s time step at selected prediction horizons. All the values have a unit of percent (%).
Prediction Horizon
30 s 35 s 40 s 45 s
Time Step mixed 0.02 -0.15 -0.22 -0.18
2/60 s $\bm{0.13}$ $\bm{0.17}$ 0.06 0.06
3/60 s -0.49 -0.43 -0.42 -0.13
4/60 s -0.14 0.05 $\bm{0.12}$0.13
5/60 s -0.33 -0.38 0.01 $\bm{0.34}$
V-D2 Robustness to parameter initial values
We conducted experiments to investigate the model’s robustness to the initial values of the physical parameters in the first phase of training. The experiments employed data from six different spiral motion configurations with $\Delta_{\overline{\bm{r}}_{x}}=2\text{cm}$. Each test involves perturbing $\bm{\eta}_{K_{i}}$ where $i=1,2,3$ with four multiplicative levels: $0.0$, $0.5$, $1.5$, and $2.0$, representing -100%, -50%, +50%, and +100% perturbation, respectively. This results in eighteen models trained at each perturbation level across the six configurations. The average losses on the test trajectories are detailed in Table IV, which also includes the losses of models without perturbations, labeled as $1.0$ for comparison. Comparing these variations to the original parameters, we observed a substantial increase in the loss of the first-principle model. Specifically, at a disturbance of $200\%$, the loss growth rate of the first-principle model was approximately 21 times greater than that of the ABNODE model, demonstrating ABNODE’s superior robust performance to inaccurate initial parameter values.
TABLE IV: Average loss comparison between various levels of multiplicative perturbation in the initial values of physical parameters using first principle and ABNODE models ($10^{-2}$).
multiplicative levels
0.0 0.5 1.0 1.5 2.0
First-principle 3.73 1.30 0.38 1.46 3.46
ABNODE 0.32 0.30 0.15 0.20 0.21
@@ VI Conclusion
This letter proposed the ABNODE model that consisted of a first-principle module with physical parameter auto-tuning and a neural network module for modeling residual dynamics in robotic blimps. Extensive spiraling experiments were conducted, the results of which demonstrated improved modeling accuracy and generalizability compared to three baseline models including the first-principle, SINDYc, and BNODE.
In future work, we plan to extend the application scope of the ABNODE model to accommodate a wider range of motions executed by robotic blimps, e.g., large attack angles, stall maneuvers, and high velocities. Additionally, we will investigate feedback control designs built upon the ABNODE model for agile and complex flight tasks by adding environmental interference in real-world application.